### v1

@@ Abstract
This paper offers educational insight into the Dirac equation, examining its historical context and contrasting it with the earlier Schrödinger and Klein-Gordon (KG) equations. The comparison highlights their Lorentz transformation symmetry and potential probabilistic interpretations. We explicitly solve the free-particle dynamics in Dirac’s model, revealing the emergence of negative-energy solutions. This discussion examines the Dirac Sea Hypothesis and explores the solutions’ inherent helicity. Additionally, we demonstrate how the Dirac equation accounts for spin and derive the Pauli equation in the non-relativistic limit. The Foldy-Wouthuysen transformation reveals how the equation incorporates spin-orbit interaction and other relativistic effects, ultimately leading to the fine structure of hydrogen. A section on relativistic covariant notation is included to emphasize the invariance of the Dirac equation, along with more refined formulations of both the KG and Dirac equations. Designed for undergraduate students interested in the Dirac equation, this resource provides a historical perspective without being purely theoretical. Our approach underscores the significance of a pedagogical method that combines historical and comparative elements to profoundly understand the role of the Dirac equation in modern physics.
@@ I Introduction
The Dirac equation, conceived by Paul Dirac in 1928 [1, 2], is one of the cornerstones of contemporary physics, merging the principles of special relativity and quantum mechanics. This equation not only predicts antiparticles but also contains the Pauli equation in the non-relativistic limit, consequently incorporating the concept of spin. Revered as a milestone in scientific history [3], it is also distinguished by its beauty, as highlighted by Wilczek [4]: “Of all the equations of physics, perhaps the most magical is the Dirac equation”. Moreover, the applications of Dirac theory span various domains, including Quantum Field Theory [5, 6], Material Science [7, 8], and Cosmology [9, 10].
Beyond its mathematical and physical significance, the equation garners attention outside academia, prominently featured in scientific dissemination materials [11, 12]. This widespread coverage underscores its emblematic importance within modern physics. However, despite its stature, Dirac’s model often remains unexplored in undergraduate studies. Hence, our proposal aims to bridge this gap by presenting a didactic introduction to the equation, not claiming to make an original contribution to Relativistic Quantum Mechanics or to the History of Physics while doing so.
@@
This work is intended for undergraduate students who are fascinated by the subject, but are not content with a purely conceptual approach. Instead, we utilize mathematical tools commonly taught in the early years of graduation, ensuring that the article remains accessible without posing significant challenges. We instructively present key concepts of quantum mechanics to the less familiar reader, aiming to do so smoothly and contextually to avoid boring the more experienced reader. For instances requiring more formal computation, we strive to clarify the steps explicitly, accompanied by further recommended readings to aid comprehension. Consciously, we have chosen not to use the Dirac-Braket notation, as it is typically not employed in the initial encounters with quantum mechanics. Exceptions are made in the appendix A, where the notation significantly facilitates the discussion regarding the dimension of the Dirac matrices, and in Subsection VI.2, where detailed calculations would be quite extensive.
Furthermore, we seek to fill a gap in references that address the incongruity between the Schrödinger equation and the relativity [3, 15, 16]. Surprisingly, even Relativistic Quantum Mechanics books [17, 18] may lack this discussion, usually starting the detailed analysis from the Klein-Gordon (KG) equation onward. To tackle this issue, we choose to demonstrate the equation’s non-invariance using a simple one-dimensional Lorentz transformation, a key principle from Relativity Theory, explained as in introductory Modern Physics classes [19]. We carry out this invariance test using the same methodology as the KG equation so that the contrast between the models is favored. This direct comparison between the wave equations is intensified so that the reader can perceive the distinctive characteristics of each equation. Reference [20] is an inspiration in this regard, as it draws an analogous comparison between the models but is restricted to the more complex case of atoms resembling hydrogen.
@@ II Historical Preamble
On 21 September 1925, Werner Heisenberg published an exploratory article titled Quantum-theoretical re-interpretation of kinematic and mechanical relations [23] in response to discrepancies between experimental observations and theoretical predictions regarding the spectral lines of the hydrogen emission spectrum [24, 208]. In this work, a significant transition was signaled, moving away from the Old Quantum Mechanics period. This era included major breakthroughs in quantum physics between 1900 and 1925, including the de Broglie hypothesis, the Pauli exclusion principle, and the Bohr-Sommerfeld atom model [25, 28]. Despite successes, the Old Quantum Mechanics exhibited methodological gaps, relying on an amalgamation of hypotheses, principles, and theorems [24]. While Bohr endeavored to reconcile spectral line discrepancies through his correspondence principle, Heisenberg diverged from such approaches and classical methods, choosing instead to ground his theory on measurable observables. Fundamentally, he departed from non observable quantities within the quantum scope. For example, instead of using the position of an electron traversing an orbit, one should utilize the probability amplitude $X_{nm}$ of the particle undergoing a transition from state $n$ to state $m$ [21]. Based on this reasoning, he formulates his well-known multiplication rule, by which two observables, such as $A_{mk}$ and $B_{kn}$, are multiplied to form a new observable $(AB)_{mn}$, which represents a combination of the original observables
@@
In Heisenberg’s algebraic framework, Born identified matrix properties, a subject rarely utilized in physics until then [24, 218]. Consequently, one would infer that Heisenberg’s observables, representing physical quantities in quantum mechanics, took the form of matrices. Born, along with Jordan, devised what would become known as matrix mechanics in their 1925 publication, "On Quantum Mechanics" [26], received merely six days after the publication of Heisenberg’s article in the same scientific journal Zeitschrift für Physik. By applying the Heisenberg rule, they successfully derived the canonical commutation relation in matrix form [21]
@@
where $h$ is Planck’s constant, $\mathbbm{1}_{n}$ is the unity matrix in dimensions $n\times n$ and $X_{nk}$ and $P_{kn}$ are matrices or observables. This asserts that it is impossible for states to measure both a definite position and momentum simultaneously, a realization that is one of the foundations of quantum mechanics. In November 1926, a follow-up to the initial paper [27], this time co-authored by Heisenberg, was published. This publication further established the principles of matrix mechanics with greater thoroughness and detail, and Born famously called it the "three-man paper".
Dirac, in turn, noticed an initially perplexing relation in Heisenberg’s article that, weeks later, would lead to his first publication in quantum mechanics: the non-commutation between certain pairs of physical quantities [16]. This relationship underwent a period of contemplation before Dirac, on a Sunday in October 1925, linked it to the Poisson brackets [28, 28], which are crucial components of Hamiltonian formalism in Classical Mechanics. In December 1925, The Fundamental Equations of Quantum Mechanics [29] is published, wherein, without using Heisenberg products or matrix properties, it was established the commutation relation
$[x,p]=i\hbar,$ (3)
@@ III Schrödinger equation
In the first semester of 1926, while Dirac was pursuing his Ph.D., Erwin Schrödinger published four works on wave mechanics [33, 34, 35, 36], all titled Quantization as an eigenvalue problem. Schrödinger’s initial goal was to derive an equation for de Broglie’s matter waves and, to this end, he used Hamilton’s optical-mechanics analogy [37, 24]. He proposed that classical mechanics represents a specific instance within the broader scope of mechanics, encompassing microscopic scenarios, a parallel to geometric optics within the wider domain of optics. Consequently, classical equations were as ineffective for quantum problems as geometric optics for wavelike problems. Therefore, it was necessary to establish wave mechanics akin to physical optics, with the wave equation replacing classical equations of motion. In hindsight, this approach reflects Dirac’s convictions about the connections of quantum mechanics with the relations in classical physics, as discussed in the preceding section.
The derivation of the Schrödinger equation presented here emphasizes some important aspects regarding the quantization framework, which will be revisited throughout the reading. We commence with the classical relationship between energy $E$ and momentum, $\boldsymbol{p}$
@@
$\boldsymbol{p}$ $\rightarrow-i\hbar\boldsymbol{\nabla},$ (7)
The relations above are a clear example of canonical quantization. Replacements (6) and (7) symbolize an important step in quantum mechanics: the transition of quantities to operators, which are specific kinds of linear transformation present in linear algebra. Replacing the aforementioned operators in (5) we obtain
@@
By employing this method, Schrödinger notably succeeded in deriving the energy levels that correspond to Bohr’s non-relativistic model for the Hydrogen atom. Schödinger [36] had an electrodynamic interpretation of his wave equation, assuming that the square modulus of $\psi\left(\mathbf{r},t\right)$ was a function of the charge distribution. However, in June 1926, Born introduced his probabilistic interpretation [39, 16, 260], considering the square modulus of $\psi\left(\mathbf{r},t\right)$ as a probability density
@@
Figure 2: Schematic representation of $\psi(x)$ and $|\psi(x)|^{2}$ for the scenario of a particle in a box of length L, a simple solution to the Schrödinger equation. Adapted from Ref. [19, 233].
@@
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}+\boldsymbol{\nabla}\cdot{\mathbf{j}(\mathbf{r},t)}=0,$ (15)
@@
where we leave the position and time variables of $\psi\left(\mathbf{r},t\right)$ implicit.
Within the frameworks of Electromagnetism and Fluid Dynamics, Eq. (15) represents the principle of charge conservation and the conservation of fluid mass, respectively. Concerning the Schrödinger equation, it signifies conservation of probability. It is natural to wonder how wave mechanics fits in with matrix mechanics. In his third article [35], Schrödinger established that wave mechanics implies some concepts from matrix mechanics. John von Neumann, only in 1929, formally demonstrated the equivalence between matrix mechanics and wave mechanics, based on a functional analysis theorem [24, 273].
An essential solution of the Schrödinger equation pertinent to this study involves a free particle. In this scenario, the differential equation allows for a solution resembling a plane wave, akin to the solution found in Maxwell’s equations within a vacuum
$\boldsymbol{\nabla}^{2}\mathbf{E}(\mathbf{r},t)-\frac{1}{c^{2}}\frac{\partial^{2}\mathbf{E}(\mathbf{r},t)}{\partial^{2}t}=0.$ (17)
If we consider the one-dimensional and time-independent case, with the associated particle confined in a box of length $L$, the Schrödinger equation admits sinusoidal solutions $\psi(x)$, as illustrated in Fig. 2, with positions $A$, $B$, and $C$ taken as examples. According to the magnitude of the squared modulus (112), it can be observed that the decreasing order of probability for these positions is $A$, $C$, and $B$. The reader more familiar with quantum mechanics recognizes that, due to the probabilistic interpretation, the integration of $\rho(\mathbf{r},t)$ over the length of the box (from $x=0$ to $x=L$) yields unity. This procedure, commonly found in textbook problems, usually seeks to identify the normalization constant of the specified wave function. The energy eigenvalues for this case are obtained through the equation
@@
from where we extract what seems to be a prerequisite, but later we will discover that it is not: positive definite energy eigenvalues.
The scientific community received the Schrödinger equation with enthusiasm, and prominent personalities such as Fritz London, Charles Darwin, and Enrico Fermi were among those who welcomed it. Its emergence was almost reassuring in the context of the departure from classical methods advocated by the newer generation of physicists [4, 110]. At the Munich Conference in the summer of 1926, most participants aligned themselves with wave mechanics to the detriment of matrix mechanics [32].
On the other hand, Dirac had an initial aversion to Schrödinger’s equation [16, 287], since he had already developed his own methodology to address quantum problems involving the aforementioned q-numbers. However, he gradually accepted the Schrödinger equation as a complement to his theory and applied wave mechanics to problems with many particles [40]; in the same paper, he introduced his transformation theory. One characteristic of this theory is that it generalizes Born’s probabilistic interpretation of the Schrödinger equation [41].
The merits of the Schrödinger wave equation are vast and are widely covered in the literature [42, 43, 44, 45]. Planck would say that “plays the same role in modern physics as the equations established by Newton, Lagrange, and Hamilton in classical mechanics” [46, 30]. However, for the purposes of this article, we will address its limitations.
Feynman asserted that the Schrödinger equation is "capable of explaining all atomic phenomena, except those involving magnetism and relativity" [44, 439]. This prompts the question: why does the equation not conform to relativity? A physical model is considered compatible with relativity when it remains invariant under Lorentz transformation, a linear transformation that converts coordinates from one reference frame to another, moving at a constant speed relative to the first, replacing the Galilean transformation of Newtonian physics. This transformation serves as a tool to assess whether a given equation aligns with the principles of relativity. A notable example of invariance is found in the solution of Maxwell’s equations for a vacuum (17), where the orders of spatial (Laplacian) and temporal derivatives are equal. On the other hand, equation (10), due to the different orders of derivatives, intuitively suggests a lack of “equality” between space and time as proposed by Einstein’s theory; this is the brief explanation provided by the majority of textbooks [3, 15, 16].
@@
Figure 3: The inertial frames $\mathcal{O}$ and $\mathcal{O}^{\prime}$. Adapted from Ref. [47].
@@
[…]
@@
[…]
@@
[…]
@@
$=\gamma{{}^{2}}\left(v{{}^{2}}\frac{\partial{{}^{2}}}{\partial x^{\prime}{{}^{2}}}-2v\frac{\partial}{\partial x^{\prime}\partial t^{\prime}}+\frac{\partial{{}^{2}}}{\partial t^{\prime}{{}^{2}}}\right).$ (26)
@@
[…]
@@
Initially, Schrödinger himself attempted to derive a valid relativistic equation, but was unsuccessful in calculating the energy eigenvalues [24, 258] according to the experimentally successful Sommerfeld energy levels, derived from the quantization of the relativistic Bohr atom. The Sommerfeld expression [50] for the energy eigenvalues of the hydrogen atom is
@@
In Eq. (35), Schrödinger obtained $n-k+1/2$ and $k-1/2$ instead of $n-k$ and $k$, respectively [3]. However, it is known that this error is due to the non incorporation of electron spin into the equation [24, 258]. Schrödinger eventually informed Dirac that he had developed a relativistic wave equation, but it was unable to reproduce the Sommerfeld formula. In particular, Dirac displayed a noteworthy perspective, stating that the colleague should have maintained confidence in his “beautiful relativistic theory”, even if it was inconsistent with accurate experimental data [28]. A more in-depth discussion of this aspect of Schrödinger’s theory can be found in [51].

### v2

@@ Abstract
This paper offers educational insight into the Dirac equation, examining its historical context and contrasting it with the earlier Schrödinger and Klein-Gordon equations. The comparison highlights their Lorentz transformation symmetry and potential probabilistic interpretations. We explicitly solve the free particle dynamics in Dirac’s model, revealing the emergence of solutions with negative energy. In this discussion, we examine the Dirac Sea Hypothesis and explore the inherent helicity of the solutions. Additionally, we demonstrate how the Dirac equation accounts for spin and derive the Pauli equation in the non-relativistic limit. Through the Foldy-Wouthuysen transformation, we reveal how the equation incorporates spin-orbit interaction and other relativistic effects, ultimately leading to the hydrogen fine structure. A section on relativistic covariant notation is included to emphasize the invariance of the Dirac equation, along with more refined formulations of both the KG and Dirac equations. Designed for undergraduate students with an interest in the Dirac equation, this resource provides a historical perspective without being purely theoretical. Our approach underscores the significance of a pedagogical method that combines historical and comparative elements to profoundly understand the role of the Dirac equation in modern physics.
Eur. J. Phys. 46 053001 (2025); doi: 10.1088/1361-6404/adf9a5
@@ I Introduction
The Dirac equation, conceived by Paul Dirac in 1928 [1, 2], is one of the cornerstones of contemporary physics, merging the principles of Special Relativity and Quantum Mechanics. This equation not only predicts antiparticles but also contains the Pauli equation in the non-relativistic limit, consequently incorporating the concept of spin. Revered as a milestone in scientific history [3], it is also distinguished by its beauty, as highlighted by Wilczek [4]: “Of all the equations of physics, perhaps the most magical is the Dirac equation”. Moreover, the applications of Dirac theory span various domains, including Quantum Field Theory [5, 6], Materials Science [7, 8], and Cosmology [9, 10].
Beyond its mathematical and physical significance, the equation garners attention outside academia, prominently featured in scientific popularization materials [11]. This widespread coverage underscores its emblematic importance within modern physics. However, despite its stature, Dirac’s model often remains unexplored in undergraduate studies. Hence, our proposal aims to bridge this gap by presenting a didactic introduction to the equation, not claiming to make an original contribution to Relativistic Quantum Mechanics or to the History of Physics while doing so.
@@
This work is intended for undergraduate students who are fascinated by the subject, but are not content with a purely conceptual approach. Instead, we utilize mathematical tools commonly taught in the early years of undergraduate education, ensuring that the article remains accessible without posing significant challenges. Although we aim to introduce the concepts progressively, a prior familiarity with non-relativistic Quantum Mechanics is recommended, particularly for subsections VI.1 and VI.2. For instances requiring more formal computation, we strive to clarify the steps explicitly, accompanied by further recommended readings to aid comprehension. Consciously, we have chosen not to use the Dirac-Braket notation, as it is typically not employed in the initial encounters with Quantum Mechanics. Exceptions are made in the appendix A, where the notation significantly facilitates the discussion regarding the dimension of the Dirac matrices, and in subsection VI.2, where detailed calculations would be quite extensive.
Furthermore, we seek to fill a gap in references that address the incongruity between the Schrödinger equation and Relativity [3, 14, 15]. Surprisingly, even Relativistic Quantum Mechanics books [16, 17] may lack this discussion, usually starting the detailed analysis from the Klein-Gordon (KG) equation onward. To tackle this issue, we choose to demonstrate the equation’s non-invariance using a simple one-dimensional Lorentz transformation, a key principle from Relativity Theory, explained as in introductory Modern Physics classes [18]. We carry out this invariance test using the same methodology as the KG equation so that the contrast between the models is favored. This direct comparison between the wave equations is intensified so that the reader can perceive the distinctive characteristics of each equation. Reference [19] is an inspiration in this regard, as it draws an analogous comparison between the models but is restricted to the more complex case of atoms resembling hydrogen.
@@
This paper is organized as follows. Sec. II provides a historical prelude to wave mechanics, with emphasis on the development of matrix mechanics. In Sec. III, we introduce the Schrödinger equation and offer an overview of its key features. Sec. IV presents the KG equation, focusing on its relativistic foundations and comparing it to the Schrödinger equation. Sec. V examines how Pauli introduced spin into the description of electromagnetic interactions within the Schrödinger equation framework. The Dirac equation is discussed in Sec. VI. Subsections VI.1 and VI.2 present two approaches to derive its non-relativistic limit, while section VII briefly introduces the covariant relativistic notation. Finally, Sec. VIII summarizes the main points developed throughout the paper. The appendix A discusses the dimension of the Dirac matrices.
@@ II Historical Preamble
On 21 September 1925, Werner Heisenberg published an exploratory article titled Quantum-theoretical re-interpretation of kinematic and mechanical relations [22] in response to discrepancies between experimental observations and theoretical predictions regarding the spectral lines of the hydrogen emission spectrum [23]. In this work, a significant transition was signaled, moving away from the Old Quantum Mechanics period. This era included major breakthroughs in quantum physics between 1900 and 1925, including the de Broglie hypothesis, the Pauli exclusion principle, and the Bohr-Sommerfeld atom model [24]. Despite successes, the Old Quantum Mechanics exhibited methodological gaps, relying on an amalgamation of hypotheses, principles, and theorems [23]. While Bohr endeavored to reconcile spectral line discrepancies through his correspondence principle, Heisenberg diverged from such approaches and classical methods, choosing instead to ground his theory on measurable observables. Fundamentally, he departed from non observable quantities within the quantum scope. For example, instead of using the position of an electron traversing an orbit, one should utilize the probability amplitude $X_{nm}$ of the particle undergoing a transition from state $n$ to state $m$ [20]. Based on this reasoning, he formulates his well-known multiplication rule, by which two observables, such as $A_{mk}$ and $B_{kn}$, are multiplied to form a new observable $(AB)_{mn}$, which represents a combination of the original observables
@@
In Heisenberg’s algebraic framework, Born identified matrix properties, a subject rarely utilized in physics until then [23]. Consequently, one would infer that Heisenberg’s observables, representing physical quantities in Quantum Mechanics, took the form of matrices. Born, along with Jordan, devised what would become known as matrix mechanics in their 1925 publication, "On Quantum Mechanics" [25], received merely six days after the publication of Heisenberg’s article in the same scientific journal Zeitschrift für Physik. By applying the Heisenberg rule, they successfully derived the canonical commutation relation in matrix form [20]
@@
where $h$ is Planck’s constant, $\mathbbm{1}_{n}$ is the unity matrix in dimensions $n\times n$ and $X_{nk}$ and $P_{kn}$ are matrices or observables. This asserts that it is impossible for states to measure both a definite position and momentum simultaneously, a realization that is one of the foundations of Quantum Mechanics. In November 1926, a follow-up to the initial paper [26], this time co-authored by Heisenberg, was published. This publication further established the principles of matrix mechanics with greater thoroughness and detail, and Born famously called it the “three-man paper” [23].
Dirac, in turn, noticed an initially perplexing relation in Heisenberg’s article that, weeks later, would lead to his first publication in Quantum Mechanics: the non-commutation between certain pairs of physical quantities [15]. This relationship underwent a period of contemplation before Dirac, on a Sunday in October 1925, linked it to the Poisson brackets [27], which are crucial components of Hamiltonian formalism in Classical Mechanics. In December 1925, The Fundamental Equations of Quantum Mechanics [28] is published, wherein, without using Heisenberg products or matrix properties, it was established the commutation relation
$x,p]=i\hbar,$ (3)
@@ III Schrödinger equation
In the first semester of 1926, while Dirac was pursuing his Ph.D., Erwin Schrödinger published four works on wave mechanics [32, 33, 34, 35], all titled Quantization as an eigenvalue problem. Schrödinger’s initial goal was to derive an equation for de Broglie’s matter waves and, to this end, he used Hamilton’s optical-mechanics analogy [36, 23]. He proposed that classical mechanics represents a specific instance within the broader scope of mechanics, encompassing microscopic scenarios, a parallel to geometric optics within the wider domain of optics. Consequently, classical equations were as ineffective for quantum problems as geometric optics was for wavelike problems. Therefore, it was necessary to establish wave mechanics akin to physical optics, with the wave equation replacing classical equations of motion. In hindsight, this approach reflects Dirac’s convictions about the connections of Quantum Mechanics with the relations in classical physics, as discussed in the preceding section.
The derivation of the Schrödinger equation presented here emphasizes some important aspects regarding the quantization framework, which will be revisited throughout the reading. We commence with the classical relationship between energy $E$ and momentum, $\mathbf{p}$
@@
$\mathbf{p}$ $\rightarrow-i\hbar\boldsymbol{\nabla},$ (7)
The relations above are a clear example of canonical quantization. Their specific form stems from Hamilton’s optical-mechanical analogy, in which the motion of a particle in a conservative force field obeys the same formal laws as the propagation of light through a refractive medium [23, 36]. In this context, they can be seen as expressing the connection between Hamilton’s equations for a particle and the corresponding wave equation. Additionally, the emergence of the quantum constant $\hbar$ is justified by its dimensional equivalence to that of action [32]. The presence of the imaginary unit arises from a separation of variables of the form $\psi(\mathbf{r},t)=u(\mathbf{r})\exp(-iEt/\hbar)$ in the derivation of the time-dependent Schödinger equation [35].
Replacements (6) and (7) symbolize an important step in Quantum Mechanics: the transition of quantities to operators, which are specific kinds of linear transformation present in linear algebra. Replacing the aforementioned operators in (5) we obtain
@@
With the equation now introduced, we proceed to outline several important aspects and merits of the Schrödinger equation. We begin by discussing the motivation behind the titles of Schrödinger’s works [32, 33, 34, 35], which present Quantum Mechanics as an eigenvalue problem.
@@
By employing this method, Schrödinger notably succeeded in deriving the energy levels that correspond to Bohr’s non-relativistic model for the Hydrogen atom [32, 23].
Having established the utility of the wave function, we now turn our attention to the nature of $\psi\left(\mathbf{r},t\right)$. Schödinger [35] had an electrodynamic interpretation of his wave equation, assuming that the square modulus of $\psi\left(\mathbf{r},t\right)$ was a function of the charge distribution. However, in June 1926, Born introduced his probabilistic interpretation [38, 15], considering the square modulus of $\psi\left(\mathbf{r},t\right)$ as a probability density
@@
Figure 2: Schematic representation of $\psi(x)$ and $|\psi(x)|^{2}$ for the scenario of a particle in a box of length L, a simple solution to the Schrödinger equation.
@@
$\frac{\partial}{\partial t}\rho(\mathbf{r},t)+\boldsymbol{\nabla}\cdot{\mathbf{j}(\mathbf{r},t)}=0,$ (15)
@@
where we leave the position and time variables of $\psi\left(\mathbf{r},t\right)$ implicit. Within the frameworks of Electromagnetism and Fluid Dynamics, Eq. (15) represents the principle of charge conservation and the conservation of fluid mass, respectively. Concerning the Schrödinger equation, it signifies conservation of probability.
It is natural to wonder how wave mechanics fits in with matrix mechanics. In his third article [34], Schrödinger established that wave mechanics implies some concepts from matrix mechanics. John von Neumann, only in 1929, formally demonstrated the equivalence between matrix mechanics and wave mechanics, based on a functional analysis theorem [23].
Next, we focus on a essential application of the Schrödinger equation: the case of a free particle. In this scenario, we set $V\left(\mathbf{r},t\right)=0$ in Eq. (10) and rearrange the terms, solving for an eigenfunction $\Psi\left(\mathbf{r},t\right)$, such that
$\frac{\hbar^{2}}{2m}\boldsymbol{\nabla}^{2}\Psi\left(\mathbf{r},t\right)+i\hbar\frac{\partial}{\partial t}\Psi\left(\mathbf{r},t\right)=0,$ (17)
which is akin to the plane wave solution found in Maxwell’s equations within a vacuum
$\boldsymbol{\nabla}^{2}\mathbf{E}(\mathbf{r},t)-\frac{1}{c^{2}}\frac{\partial^{2}}{\partial^{2}t}\mathbf{E}(\mathbf{r},t)=0.$ (18)
A notable distinction is the presence of the second-order time derivative in the latter, which is absent in the former. This aspect will be addressed in a subsequent discussion.
Still within the context of the free particle, we consider the particle-in-a-box case, a concrete example of energy quantization resulting from boundary conditions. We consider the one-dimensional and time-independent case, with the associated particle confined in a box of length $L$, the Schrödinger equation admits sinusoidal solutions $\Psi(x)$, as illustrated in Fig. 2, with positions $A$, $B$, and $C$ taken as examples. According to the magnitude of the squared modulus (113), it can be observed that the decreasing order of probability for these positions is $A$, $C$, and $B$. The reader more familiar with Quantum Mechanics recognizes that, due to the probabilistic interpretation, the integration of $\rho(\mathbf{r},t)$ over the length of the box (from $x=0$ to $x=L$) yields unity. This procedure, commonly found in textbook problems, usually seeks to identify the normalization constant of the specified wave function. The energy eigenvalues for this case are obtained through the equation
@@
from where we extract what seems to be a prerequisite, but later we will discover that it is not: positive definite energy eigenvalues. We further note that a given state $\psi(x)$ can be expressed as a linear combination of the eigenfunctions $\Psi(x)$, as a consequence of the linearity of the Schrödinger equation.
The scientific community received the Schrödinger equation with enthusiasm, and prominent personalities such as Fritz London, Charles Darwin, and Enrico Fermi were among those who welcomed it. Its emergence was almost reassuring in the context of the departure from classical methods advocated by the newer generation of physicists [4]. At the Munich Conference in the summer of 1926, most participants aligned themselves with wave mechanics to the detriment of matrix mechanics [31].
On the other hand, Dirac had an initial aversion to Schrödinger’s equation [15], since he had already developed his own methodology to address quantum problems involving the aforementioned q-numbers. However, he gradually accepted the Schrödinger equation as a complement to his theory and applied wave mechanics to problems with many particles [39]; in the same paper, he introduced his transformation theory. One characteristic of this theory is that it generalizes Born’s probabilistic interpretation of the Schrödinger equation [40].
The merits of the Schrödinger wave equation are vast and are widely covered in the literature [41, 42, 43, 44]. Planck would say that it “plays the same role in modern physics as the equations established by Newton, Lagrange, and Hamilton in classical mechanics” [45]. However, for the purposes of this article, we will address its limitations.
Feynman asserted that the Schrödinger equation is "capable of explaining all atomic phenomena, except those involving magnetism and relativity" [43]. This prompts the question: why does the equation not conform to Relativity? A physical model is considered compatible with Relativity when it remains invariant under Lorentz transformation, a linear transformation that converts coordinates from one reference frame to another, moving at a constant speed relative to the first, replacing the Galilean transformation of Newtonian physics. This transformation serves as a tool to assess whether a given equation aligns with the principles of Relativity. A notable example of invariance is found in the solution of Maxwell’s equations for a vacuum (18), where the orders of spatial (Laplacian) and temporal derivatives are equal. On the other hand, equation (10), due to the different orders of derivatives, intuitively suggests a lack of “equality” between space and time as proposed by Einstein’s theory; this is the brief explanation provided by the majority of textbooks [3, 14, 15].
@@
Figure 3: The inertial frames $\mathcal{O}$ and $\mathcal{O}^{\prime}$.
@@
$={}$ $\gamma\frac{\partial}{\partial x^{\prime}}-\gamma\frac{v}{c{{}^{2}}}\frac{\partial}{\partial t^{\prime}}.$ (24)
@@
$={}$ $\gamma\frac{\partial}{\partial x^{\prime}}\left(\gamma\frac{\partial}{\partial x^{\prime}}-\gamma\frac{v}{c{{}^{2}}}\frac{\partial}{\partial t^{\prime}}\right)-\gamma\frac{v}{c{{}^{2}}}\frac{\partial}{\partial t^{\prime}}\left(\gamma\frac{\partial}{\partial x^{\prime}}-\gamma\frac{v}{c{{}^{2}}}\frac{\partial}{\partial t^{\prime}}\right),$
$={}$ $\gamma^{2}\left(\frac{\partial^{2}}{\partial x^{\prime 2}}-2\frac{v}{c{{}^{2}}}\frac{\partial^{2}}{\partial x^{\prime}\partial t^{\prime}}+\frac{v^{2}}{c^{4}}\frac{\partial^{2}}{\partial t^{\prime 2}}\right).$ (25)
@@
$={}$ $-\gamma v\frac{\partial}{\partial x^{\prime}}+\gamma\frac{\partial}{\partial t^{\prime}}.$ (26)
@@
[…]
@@
$={}$ $-v\frac{\partial}{\partial x^{\prime}}+\frac{\partial}{\partial t^{\prime}}.$ (33)
@@
Initially, Schrödinger himself attempted to derive a valid relativistic equation, but was unsuccessful in calculating the energy eigenvalues [23] according to the experimentally successful Sommerfeld energy levels, derived from the quantization of the relativistic Bohr atom. The Sommerfeld expression [48] for the energy eigenvalues of the hydrogen atom is
@@
In Eq. (36), Schrödinger obtained nk+1/2n-k+1/2 and k1/2k-1/2 instead of nkn-k and $k$, respectively [3]. However, it is known that this error is due to the non incorporation of electron spin into the equation [23]. Schrödinger eventually informed Dirac that he had developed a relativistic wave equation, but it was unable to reproduce the Sommerfeld formula. In particular, Dirac displayed a noteworthy perspective, stating that the colleague should have maintained confidence in his “beautiful relativistic theory”, even if it was inconsistent with accurate experimental data [27]. A more in-depth discussion of this aspect of Schrödinger’s theory can be found in [49].